Circular No. 5684

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 QB_{1}

The following observations have been reported since the October batch of Minor Planet Circulars was prepared (cf. IAUC 5633):

1992 UT  α_{2000}  δ_{2000}  R Observer
Sept. 24.37682 23^{h}59^{m}22.^{s}62 −0^{o}02′50′′.3  Rohde
27.37038 23 59 08.96 −0 04 17.0  _{”}
Oct. 26.42932 23 57 07.30 −0 17 10.6  de Costa
26.62049 23 57 06.59 −0 17 15.9  _{”}
Nov. 15.15329 23 56 08.51 −0 23 22.4  Monet
16.16073 23 56 06.26 −0 23 35.9  _{”}
23.05860 23 55 53.08 −0 24 57.5 23.5 Storm
Dec. 25.07396 23 55 47.23 −0 25 18.4  Hainaut

J. R. Rohde, D. Pascu and A. K. B. Monet (U.S. Naval Observatory, Flagstaff Station). 1.55-m astrometric reflector + CCD. Measurer J. A. DeYoung.

G. de Costa and K. Taylor (Anglo-Australian Observatory). Measurer R. H.

McNaught. J. Storm, A. Smette and O. Hainaut (European Southern Observatory).

3.5-m New Technology Telescope + EMMI. Measurer Hainaut.

The following improved orbital elements, by B. G. Marsden, satisfy the 20 available observations Aug. 30-Dec. 25 with mean residual 0′′.45:

Epoch = 1993 Jan. 13.0 TT

T = 2023 Aug. 10.4380 TT  ω = 46.^{o}3760
e = 0.106879  Ω = 359.4152 }2000.0
q = 39.641983 AU  i = 2.2221
a = 44.385886 AU n^{o} = 0.0033330 P = 295.7 years

The rather extreme precision is provided for the purpose of consistency, not accuracy. Considerable uncertainty should still be attached to most of the elements. It does seem, however, that the object is in a large, lowinclination orbit and relatively close to its perihelion. It is quite likely that q is significantly smaller and e significantly larger than shown above, but it is unlikely that the object could be Neptune-crossing by more than a very small amount.

1992 December 30 Brian G. Marsden